# Financial Risk Relation Identification through Dual-view Adaptation

### v1

@@ Abstract
A multitude of interconnected risk events—ranging from regulatory changes to geopolitical tensions—can trigger ripple effects across firms. Identifying inter-firm risk relations is thus crucial for applications like portfolio management and investment strategy. Traditionally, such assessments rely on expert judgment and manual analysis, which are, however, subjective, labor-intensive, and difficult to scale. To address this, we propose a systematic method for extracting inter-firm risk relations using Form 10-K filings—authoritative, standardized financial documents—as our data source. Leveraging recent advances in natural language processing, our approach captures implicit and abstract risk connections through unsupervised fine-tuning based on chronological and lexical patterns in the filings. This enables the development of a domain-specific financial encoder with a deeper contextual understanding and introduces a quantitative risk relation score for transparency, interpretable analysis. Extensive experiments demonstrate that our method outperforms strong baselines across multiple evaluation settings.
@@ 1 Introduction
To overcome these limitations, we adopt a retrieval-based encoding framework—a foundation of modern NLP—that transforms text into dense, semantically rich vector representations. This enables efficient relation discovery, information retrieval, and downstream financial NLP applications Wang et al. (2024); Alaparthi and Mishra (2020). Specifically, we first propose an unsupervised fine-tuning strategy based on a dual-view similarity framework to adapt general-purpose encoders for the finance domain. Our approach leverages two key characteristics of Form 10-K filings: (1) standardized language and constrained vocabulary, which yield consistent lexical patterns; and (2) frequent association of risk events with date-time references, enabling temporal alignment. By modeling both lexical and chronological similarities, we construct high-quality positive training pairs that reflect semantically or temporally aligned content. This dual-view supervision guides the encoder to capture nuanced financial semantics and align similar risk disclosures across firms.
To complement this encoder, we secondly introduce a retrieval-based, interpretable scoring mechanism—the risk relation score (RRS)—to quantify inter-firm risk connections. RRS offers key advantages over traditional heuristics: it is symmetric, guarantees minimum similarity thresholds, and enhances interpretability by grounding each relation in explicitly retrieved mutual risk paragraphs (MRPs). This not only provides a robust measure of shared risk but also delivers textual evidence, improving the transparency and reliability of the model’s output. Figure 1 gives an overview of our method for identifying inter-firm risk relations. [^3] A peer-reviewed system demonstration related to this work was independently developed and published as a demo paper at NAACL 2025 by Wang et al. (2025). Although both the demo paper and the present paper employ similar methodology, the former primarily emphasizes the visualization of outcomes, whereas the latter focuses on the methodological contributions and comprehensive empirical experiments across various aspects.
@@
    We assess the standalone retrieval capabilities of our encoder using MultiHiertt Zhao et al. (2022), a financial QA benchmark built from regulatory filings, where our model consistently outperforms strong baselines.
@@ 2.1 Relation Extraction
Early relation extraction (RE) methods relied on pattern matching and manual feature engineering, such as such as the lexico-syntactic patterns introduced by Hearst (1992). With the rise of deep learning, neural models have like CNN Zeng et al. (2015) and PCNN Zeng et al. (2015) improved RE through automated feature learning. The advent of pretrained language models, particularly BERT Devlin et al. (2019), significantly advanced RE by enabling fine-tuning for contextual understanding. Approaches such as entity-aware fine-tuning Soares et al. (2019) further boosted performance. Recent surveys Diaz‑Garcia and Lopez (2024) highlighted the dominance of BERT-based techinques while recognizing the growing impact of large language models (LLMs) like T5 Raffel et al. (2020).
@@ 2.2 General-domain Encoders
Transformer-based models, led by BERT Devlin et al. (2019) and its variants such as RoBERTa Liu et al. (2019) and SpanBERT Joshi et al. (2020), transformed NLP by enabling transfer learning across tasks. In the context of retrieval, DPR Karpukhin et al. (2020) introduced supervised dense retrieval with dual encoders, while Spider Ram et al. (2021) used instruction-tuned data to enhance performance. Contriever Izacard et al. (2021) offered an unsupervised contrastive learning alternative, effectively capturing semantic similarity. More recent models, such as Jina AI’s encoder Günther et al. (2023), incorporates novel methods like Attention with Linear Biases (ALiBi) Press et al. (2021) to effectively process longer textual sequences. Embeddings in Xiao et al. (2024) further refine encoding strategies by considering context from multiple perspectives including functionality, granularity, and linguity.
@@ 2.3 Financial-domain Encoders
Domain-specific encoders have been introduced to better capture the nuances of financial language. FinBERT Araci (2019) adapted BERT by pretraining on large-scale financial corpora and has demonstrated superior performance in sentiment analysis. SEC-BERT Loukas et al. (2022) further narrowed the focus by training exclusively on U.S. SEC filings, enhancing its applicability to regulatory documents. Another variant, FinBERT-MRC Zhang and Zhang (2023) reformulated financial named entity recognition as a machine reading comprehension task, improving contextual precision. Beyond BERT-based models, Fin-E5 Tang and Yang (2025) introduced a persona-driven synthetic data strategy to support a wider range of financial embedding tasks. In parallel, proprietary models such as BloombergGPT Wu et al. (2023) showcased the potential of financial LLMs trained on exclusive datasets, though their closed nature spurred demand for open alternatives. In response, FinGPT Yang et al. (2023) offers an open-source framework focused on accessible data and democratized financial LLMs.
@@ 3.2.1 Unsupervised Training of the Encoder
We fine-tune the encoder using contrastive learning to bring similar paragraphs closer in embedding space. Given an anchor text piece $p$, a positive counterpart $p^{+}$, and a set of negatives $D^{-}$, the InfoNCEloss van den Oord et al. (2018) is defined as:
@@
This setup provides B1B-1 negative examples per anchor, enhancing the contrastive signal. While positive pairs are constructed based on the chronological and lexical similarity views described in Section 3.2.1, negative pairs are dynamically formed from the remaining batch entries. This strategy yields a robust and scalable training procedure, consistent with techniques from prior work Chen et al. (2017, 2020).
@@ 4.1 Encoder Training Details
For model training, we construct 8,500 positive paragraph pairs and 1,000 validation pairs for each of the two views: chronological and lexical similarity. We fine-tune our encoder starting from the BERT-base-uncased pretrained model [^4] https://huggingface.co/google-bert/bert-base-uncased. Training is performed using contrastive learning with a batch size of 64 and a learning rate of $2\times 10^{-5}$, optimized via Adam optimizer Kingma and Ba (2014) with a linear warmup scheduler. L2 regularization is applied to improve generalization, and training proceeds for up to 50 epochs with early stopping to mitigate overfitting. [^5] For computational resources, training was conducted on a single NVIDIA V100 GPU for approximately 8 hours.
@@ 4.3 Encoders for Comparison
  • Contriever Izacard et al. (2021): An unsupervised retrieval model trained with contrastive learning.
  • DPR Karpukhin et al. (2020): A supervised retrieval model trained on question-answer pairs.
  • FinBERT Araci (2019) and SEC-BERT Loukas et al. (2022): Two domain-specific models pretrained on financial corpora. We use the base version of SEC-BERT to ensure fairness in model size.
  • Llama-3.2 Touvron et al. (2024): A widely recognized open-source large language model (LLM). We employ the 3B variant to provide a fair comparison in terms of model size with our encoder.
@@ 4.7 Applying Risk Relations to Stock Price Movement Prediction (RQ3)
We evaluate the practical utility of our risk relation metric by applying it to a downstream financial task: stock price movement prediction. Specifically, we integrate our risk relations into the attribute-driven graph attention networks (ADGAT) Cheng and Li (2021) to enhance stock price movement prediction. The original ADGAT framework notes that using human-defined relations (e.g., sector or industry links) often introduces biases and degrades performance, as such links are static, binary, and lack contextual nuance. To address this, we replace ADGAT’s predefined relations with those derived from our method.
@@ 4.7.1 Experimental Setup
Each configuration is run 15 times with different seeds, and the averages of the top 5 results are reported. To evaluate statistical significance, we apply a Mann–Whitney U test Mann and Whitney (1947) to compare performance distributions.
@@ 4.8.2 Experimental Setup
We extract 290 queries and their 1,331 corresponding relevant paragraphs. A two-stage retrieval pipeline is adopted. In the first stage, we retrieve the top 200 documents based on cosine similarities between the query and document embeddings. In the second stage, the candidates are re-ranked using the BAAI/bge-reranker-v2-m3 Li et al. (2023); Chen et al. (2024). To evaluate our encoder, we replace the retriever component in this pipeline with ours and compare its performance against several baseline encoders.

### v2

@@ Abstract
A multitude of interconnected risk events—ranging from regulatory changes to geopolitical tensions—can trigger ripple effects across firms. Identifying inter-firm risk relations is thus crucial for applications like portfolio management and investment strategy. Traditionally, such assessments rely on expert judgment and manual analysis, which are, however, subjective, labor-intensive, and difficult to scale. To address this, we propose a systematic method for extracting inter-firm risk relations using Form 10-K filings—authoritative, standardized financial documents—as our data source. Leveraging recent advances in natural language processing, our approach captures implicit and abstract risk connections through unsupervised fine-tuning based on chronological and lexical patterns in the filings. This enables the development of a domain-specific financial encoder with a deeper contextual understanding and introduces a quantitative risk relation score for transparency, interpretable analysis. Extensive experiments demonstrate that our method outperforms strong baselines across multiple evaluation settings. Our codes are available at https://github.com/cnclabs/codes.fin.relation.
@@ 1 Introduction
To overcome these limitations, we adopt a retrieval-based encoding framework—a foundation of modern NLP—that transforms text into dense, semantically rich vector representations. This enables efficient relation discovery, information retrieval, and downstream financial NLP applications wang2024bertir; alaparthi2020bert-sa. Specifically, we first propose an unsupervised fine-tuning strategy based on a dual-view similarity framework to adapt general-purpose encoders for the finance domain. Our approach leverages two key characteristics of Form 10-K filings: (1) standardized language and constrained vocabulary, which yield consistent lexical patterns; and (2) frequent association of risk events with date-time references, enabling temporal alignment. By modeling both lexical and chronological similarities, we construct high-quality positive training pairs that reflect semantically or temporally aligned content. This dual-view supervision guides the encoder to capture nuanced financial semantics and align similar risk disclosures across firms.
To complement this encoder, we secondly introduce a retrieval-based, interpretable scoring mechanism—the risk relation score (RRS)—to quantify inter-firm risk connections. RRS offers key advantages over traditional heuristics: it is symmetric, guarantees minimum similarity thresholds, and enhances interpretability by grounding each relation in explicitly retrieved mutual risk paragraphs (MRPs). This not only provides a robust measure of shared risk but also delivers textual evidence, improving the transparency and reliability of the model’s output. Figure 1 gives an overview of our method for identifying inter-firm risk relations. [^3] A peer-reviewed system demonstration related to this work was independently developed and published as a demo paper at NAACL 2025 by wang-etal-2025-surf. Although both the demo paper and the present paper employ similar methodology, the former primarily emphasizes the visualization of outcomes, whereas the latter focuses on the methodological contributions and comprehensive empirical experiments across various aspects.
@@
    We assess the standalone retrieval capabilities of our encoder using MultiHiertt zhao-etal-2022-multihiertt, a financial QA benchmark built from regulatory filings, where our model consistently outperforms strong baselines.
@@ 2.1 Relation Extraction
Early relation extraction (RE) methods relied on pattern matching and manual feature engineering, such as such as the lexico-syntactic patterns introduced by hearst1992hyponyms. With the rise of deep learning, neural models have like CNN zeng-etal-2015-pcnn and PCNN zeng-etal-2015-pcnn improved RE through automated feature learning. The advent of pretrained language models, particularly BERT devlin-etal-2019-bert, significantly advanced RE by enabling fine-tuning for contextual understanding. Approaches such as entity-aware fine-tuning baldini-soares-etal-2019-mtb further boosted performance. Recent surveys diaz2024re-survey highlighted the dominance of BERT-based techinques while recognizing the growing impact of large language models (LLMs) like T5 raffel2020t5.
@@ 2.2 General-domain Encoders
Transformer-based models, led by BERT devlin-etal-2019-bert and its variants such as RoBERTa liu-etal-2019-roberta and SpanBERT joshi-etal-2020-spanbert, transformed NLP by enabling transfer learning across tasks. In the context of retrieval, DPR karpukhin-etal-2020-dpr introduced supervised dense retrieval with dual encoders, while Spider ram2021spider used instruction-tuned data to enhance performance. Contriever izacard2021unsupervised offered an unsupervised contrastive learning alternative, effectively capturing semantic similarity. More recent models, such as Jina AI’s encoder gunther2023jina, incorporates novel methods like Attention with Linear Biases (ALiBi) press2021alibi to effectively process longer textual sequences. Embeddings in xiao2024cpack further refine encoding strategies by considering context from multiple perspectives including functionality, granularity, and linguity.
@@ 2.3 Financial-domain Encoders
Domain-specific encoders have been introduced to better capture the nuances of financial language. FinBERT araci2019finbert adapted BERT by pretraining on large-scale financial corpora and has demonstrated superior performance in sentiment analysis. SEC-BERT loukas2022finer further narrowed the focus by training exclusively on U.S. SEC filings, enhancing its applicability to regulatory documents. Another variant, FinBERT-MRC zhang2023finbert-mrc reformulated financial named entity recognition as a machine reading comprehension task, improving contextual precision. Beyond BERT-based models, Fin-E5 tang2025finmteb introduced a persona-driven synthetic data strategy to support a wider range of financial embedding tasks. In parallel, proprietary models such as BloombergGPT wu2023bloomberggpt showcased the potential of financial LLMs trained on exclusive datasets, though their closed nature spurred demand for open alternatives. In response, FinGPT yang2023fingpt offers an open-source framework focused on accessible data and democratized financial LLMs.
@@ 3.2.1 Unsupervised Training of the Encoder
We fine-tune the encoder using contrastive learning to bring similar paragraphs closer in embedding space. Given an anchor text piece $p$, a positive counterpart $p^{+}$, and a set of negatives $D^{-}$, the InfoNCEloss oord-etal-2018-cpc is defined as:
@@
This setup provides B1B-1 negative examples per anchor, enhancing the contrastive signal. While positive pairs are constructed based on the chronological and lexical similarity views described in Section 3.2.1, negative pairs are dynamically formed from the remaining batch entries. This strategy yields a robust and scalable training procedure, consistent with techniques from prior work chen2017sampling; chen2020simclr.
@@ 4.1 Encoder Training Details
For model training, we construct 8,500 positive paragraph pairs and 1,000 validation pairs for each of the two views: chronological and lexical similarity. We fine-tune our encoder starting from the BERT-base-uncased pretrained model [^4] https://huggingface.co/google-bert/bert-base-uncased. Training is performed using contrastive learning with a batch size of 64 and a learning rate of $2\times 10^{-5}$, optimized via Adam optimizer kingma2014adam with a linear warmup scheduler. L2 regularization is applied to improve generalization, and training proceeds for up to 50 epochs with early stopping to mitigate overfitting. [^5] For computational resources, training was conducted on a single NVIDIA V100 GPU for approximately 8 hours.
@@ 4.3 Encoders for Comparison
  • Contriever izacard2021unsupervised: An unsupervised retrieval model trained with contrastive learning.
  • DPR karpukhin-etal-2020-dpr: A supervised retrieval model trained on question-answer pairs.
  • FinBERT araci2019finbert and SEC-BERT loukas2022finer: Two domain-specific models pretrained on financial corpora. We use the base version of SEC-BERT to ensure fairness in model size.
  • Llama-3.2 touvron2024llama3: A widely recognized open-source large language model (LLM). We employ the 3B variant to provide a fair comparison in terms of model size with our encoder.
@@ 4.7 Applying Risk Relations to Stock Price Movement Prediction (RQ3)
We evaluate the practical utility of our risk relation metric by applying it to a downstream financial task: stock price movement prediction. Specifically, we integrate our risk relations into the attribute-driven graph attention networks (ADGAT) cheng-li-2021-adgat to enhance stock price movement prediction. The original ADGAT framework notes that using human-defined relations (e.g., sector or industry links) often introduces biases and degrades performance, as such links are static, binary, and lack contextual nuance. To address this, we replace ADGAT’s predefined relations with those derived from our method.
@@ 4.7.1 Experimental Setup
Each configuration is run 15 times with different seeds, and the averages of the top 5 results are reported. To evaluate statistical significance, we apply a Mann–Whitney U test mann1947u to compare performance distributions.
@@ 4.8.2 Experimental Setup
We extract 290 queries and their 1,331 corresponding relevant paragraphs. A two-stage retrieval pipeline is adopted. In the first stage, we retrieve the top 200 documents based on cosine similarities between the query and document embeddings. In the second stage, the candidates are re-ranked using the BAAI/bge-reranker-v2-m3 li2023llmretrieval; chen2024bgem3. To evaluate our encoder, we replace the retriever component in this pipeline with ours and compare its performance against several baseline encoders.